A car has to cover the distance 60 km, if half of the total time it travels with velocity 80 km/h and in rest half time its speed becomes 40 km/h, the average speed of car will be,

(a) 60 km/h         (b) 80 km/h

(c) 120 km/h       (d) 180 km/h

How do we solve this question?

Grade:12

2 Answers

Paarth Arkadi
8 years ago
The problem can be solved by two methods
Method 1 :
Vav = 80 + 40 / 2
       = 120/2
       = 60 km /h
Method 2 : 
Let the half time intervals be t
Therefore 60 = 80t + 40t
t=1/2
total time = 1/2*2 = 1hr
vav = Total distance / Total time 
      = 60 / 1
     = 60 km/h
Agam Dogra
8 years ago
Question can be asked in two ways;
1)when paritcle travel with speed vin time t/2 and speed v2 in time t/2 then Vav=v1+v2/2 (when time interval is equal).
2)when paritcle travel with speed v1 for distance x and speed v2 distance x then Vav =2v1v2/v1+v2   (when distance covered is equal).